OpenStudy (anonymous):

The cost of a ticket to an amusement park was increased by 4 dollars. The new cost is twenty-eight dollars. a. Write a one-step equation that could be solved to find the old price. Let p represent the old price. b. Solve the equation you wrote in Part a. Show all your work. c. Explain how you know that the solution you found in Part b is the correct answer.

OpenStudy (anonymous):

pleaee help me

OpenStudy (anonymous):

A) 28=p+4 B) 28-4=p+4-4 =>24=p as per statement..new price is 4 greater then old price..so it can be judge from above explanation..

OpenStudy (anonymous):

can you help me put this in a word problem

OpenStudy (anonymous):

whats that..?

OpenStudy (anonymous):

it a word problem